4. (a) Test the positive definiteness of the following matrices. 2 -1 (i) A = -1 2 0 0 29 7 (ii) A= 4 5 6 3 2 3
Answer #1

2.4 9) rin A= Nii ī no Ton det A= ? -1 í ro No. =2 (4) + (-2)-1(2) = 8-2-2 detA= 4 det A to and all principal diagonal elemen(۱) A= 9 u 5 wat 3 २ det A= 7 6 3 4 5 3 eurao = 9(1572) -8 (17-18)+7(8-15) Idet =26) deta to, and all principal diagonal elem